To whom may it concern,
I have problem to solve this questions :-
Peter and Jane together have $90. If Peter gave Jane $16, he would then have $14 less than Jane. How much does Peter has originally ?
Thank you.

Found 2 solutions by checkley71, cyyn03:
Answer by checkley71(8403)

P+J=90 & P-16=J-14
OR P=90-J OR 90-J-16=J-14 OR -2J=-90+16-14 OR -2J=-88 OR J=88/2 OR J=44 THUS
P+44=90 OR P=90-44 OR P=46
PROOF 46-16=44-14 OR 30=30

Answer by cyyn03(1)

J+P=90; (J+16)-(P-16)=14
Solve:
J+16-P+16=14
J-P=14-32
P=J+18
P=(90-P)+18
P+P=90+18
2P=108
P=108/2=54
Peter = $54
Check:
36+54=90
36+16=52; 54-16=38, so 52-38=14